Home‎ > ‎Reports‎ > ‎Crew 109‎ > ‎

day04


Crew #109 Log Book for December 20, 2011

Photos of the day





Commander's Check-In Report

Date and time: 20 December 2011, 19:30

Written by Nicky De Munster, Crew 109, Commander:

Crew Physical Status: good

Time Departed/Returned from EVA:

EVA5: 8:45 - 11:30
EVA6: 1.35 - 4:16

Brief Narrative of Field Mission Results:

EVA 5 collected core samples and tested soil strength.
EVA 6 departed to search for water by map; trail not found to get there
Preparing seeds to plant for tomorrow

EVA Data/Interpretations:

Water not found
Soil samples in different places

Engineering/Hab Maintenance:

Backpack 6 repaired by Leandro

Report Transmission Schedule:

Daily reports: CMDR check-in, CMDR report, EVA report by David, ENG report by Leandro, Science report by David, Karon, Victoria

Plans for Tomorrow:

Because there isn't any gas left in any of the jerrycans or in the main gas tank, we shall do only one EVA with the ATV's and one by foot.
We are waiting to get resupplied.
We will plant seeds in Utah soil mixed with perlite
Drinking water will be analyzed for conductivity and oxygen

Inventory: n/a

Support Requested:

Gas
Water

Miscellaneous:

Commander's Report

Date and time: 20 December 2011 20:00
Written by Nicky De Munster, Crew 109, Commander

Text:

After breakfast the morning EVA crew left for a long and distant exploration. They successfully performed a test with cone penetrometer.
In the afternoon EVA crew went looking for places where water could be found.  Much time was lost because the tracks were not identifiable.
Engineering was done on backpack 6 by Leandro.
We are concerned because the water level is very low and we still were not supplied. Same with the fuel for the ATV.
When we asked whether we could use the microscopes in cabinet 8, we got a pretty aggressive response from the science team. The whole team was very impressed. Karon get no answer to her questions.
We hope this will be clarified.
The mood in the team is excellent and the team spirit is great.

Science Report

Date: 20 December 2011, 12:50


Written by David Kutai Weiss, Crew Geologist & Executive Officer, Karon Wynne, Health & and Safety Office and Crew Biologist, Victoria Grudzinski, Crew Biologist


Biology Report:


Today EVA 5 departed to collect core samples and soil strength. During this mission, Victoria collected two moist soil samples at a peak near location K, both from underneath rocks. EVA 6 set out on a mission to search for water. We located a body of water on the map and made several attempts to get there. Unfortunately, many of the trails are washed out and we were not able to reach the body of water. Karon collected some more soil samples today in different locations. We hope to find the pond on a future EVA.

Geology Report:


After collecting 2 core samples and 6 soil strength measurements at 2 different locations yesterday, there were only 8 more target sites left to hit. Already ahead of schedule, the crew of EVA 5 set out to collect 1 core sample and 3 soil strength measurements at only 1 location. David, Leandro, and Victoria departed to site K to begin work. In transit, the crew got sidetracked in an effort to explore new trails. The upside to this detour was that David discovered some mineral samples, large amounts of red jasper and chert, types of microcrystalline quartz, as well as what the crew believes to be fluorite.


After getting their bearings, the crew departed and finally reached site K. The crew conducted 3 trials of the blowcount test with the cone penetrometer and took a core sample. The crew then took to exploring a nearby peak, where they observed large amounts of conglomerate upon reaching the summit. The crew of EVA 6 also made some interesting geological finds; Nicky collected 3 rock samples. After returning to the habitat, David was able to identify them as Calcite.

EVA 5&6 Report

Date and Time: 20 December 2011 7:30-12:50

Written by David Kutai Weiss, Crew 109 Geologist & Executive Officer

EVA 5:
Time: 8:45 am – 11:45 am
Crew Members: David, Leandro, Victoria
Site Location: East of Hab at the edge of unnamed cliff rim – site K
Transit Mode: ATV, on foot.
Objectives: General Eastern exploration. Travel to site K to perform a blowcount test with a cone penetrometer and to acquire core soil samples.

Results:

Successful traverse to site K, where morning EVA crew successfully performed a blowcount test with cone penetrometer, and acquired a core soil sample.
Victoria acquired soil samples on the nearby peak.
Morning EVA crew climbed to top of the nearby peak and had great field of view, which made for some great photos.
Leandro brought along a new map and successfully tested his new map with GPS for more accurate navigation.
Route spanned 22.4 km.

EVA 6:
Time: 13:30– 16:15
Crew Members: Nicky, Aster, Karon
Site Location: Traveled west on Sagan Street, then south on Copernicus Highway along the base of Skyline Rim. Traveled North up Old Lowell Highway until reaching Brahe Highway.
Transit Mode: ATV, on foot.
Objectives: Explore Copernicus Highway and Radio Ridge Road, and find water.

Results:

Southern stretch of Copernicus Highway tracks were not visible, so we stopped and turned back. Northern part of Radio Ridge Road tracks were not visible, so we turned back. We then searched for a pond 1 km south of Brahe highway, but did not locate it. Further mapping initiative back at the Hab showed that we drove past it, but it was hidden by the topography.
Route spanned 27 km.

EVA 5 & 6 Lessons learned:

When searching for something specifically, such as water, have a higher resolution map on hand to pinpoint its location instead of just a general area. High resolution maps with previous EVAs and coordinates mapped out combined with topographic maps and GPS seems to be a good way to navigate. Driving standing up on the ATV is quite stable.

Engineering Report

Date and time: 20 December 2011, 18:33

Written by Leandro G. Barajas, Crew 109, Chief Engineer:

Generator/Electricity:
Did you have any electricity outage? (Yes/No): NO

Fuel Status:
Diesel (Full/0.75/0.5/0.25/Empty): 0.25
Propane Tank (Full/0.75/0.5/0.25/Empty): 0.75
Gasoline Tank (Full/0.75/0.5/0.25/Empty): Empty
Notes/Comments:
All 3 ATVs fuel tanks are about 1/2 full for tomorrow's EVA however there is no gas left in any of the jerrycans or the main gas tank.
We will need to get resupplied with gas by tomorrow or we will have to stop ATV EVAs after that.

Oil:
ATV Oil Quantity (# of Quarts in storage): 1.3
Car Oil Quantity (# of Quarts in storage): N/A
Water Status:
Outside Potable Water Tank Level (inches from bottom): 12
Trailer Potable Water Tank Level (inches from bottom): 4 (it is not possible to pump more water out of it)
Hab Potable Water Tank Level (inches from bottom): 10
Potable Water Meter Reading: 47900.1
In to GreenHab Meter Reading: N/A

Notes/Comments:
CRITICAL: We need to get resupplied with water in the next 24-48 hours.

GreenHab:
Condition of plants in Tank 1: N/A
Condition of plants in Tank 2: N/A
Condition of plants in Tank 3: N/A
Greenhab Notes/Comments: N/A

Observatory:
Telescope Used? (Yes/No): NO
Observatory Notes/Comments: N/A

Transportation:

Hab Car (NOT RENTAL)
Used (yes/no): NO
Oil Checked (yes/no): YES
Tire Status: Good
Odometer at end of day: 125475.5
Notes/Comments on Hab Car:

ATV 1 (Opportunity):
Used (yes/no): YES
Oil Checked (yes/no): YES
Fuel Consumed: 2 Gal
Tire Status: Good - 4 psi
Comments:

ATV 2 (Spirit):
Used (yes/no): YES
Oil Checked (yes/no): YES
Fuel Consumed: 2 Gal
Tire Status: Good - 4 psi
Comments:

ATV 3 (Viking I):
Used (yes/no): YES
Oil Checked (yes/no): YES
Fuel Consumed: 2 Gal
Tire Status: Good - 4 psi
Comments:

Heating and Ventilation:
Thermostat setting upstairs (Farenheit): 65
Actual temperature upstairs (Farenheit): 72

Thermostat setting downstairs (Farenheit): 55
Actual temperature downstairs (Farenheit): 62

Computers/Networking Infrastructure:
Notes/comments on internet/computers?:

General Engineering Notes/Comments:

Heat Tapes Operating 24/7 (Both Lights On)

Replaced power switch in suit 5. Finish rebuild of the backpack hoses and enclosure. Suit 5 is now operational.

Fixed air hoses in suit 6. Enclosure is broken and has been temporarily fixed with duck tape. Suit 6 is now operational, however, its structural integrity and air flow are not optimal. This suit should not be used for long EVAs.